Calculate the frequency of the a allele in a popul
Answer #1

Here, the total number of individual with aa genotype = 20

So, the total frequency of aa individuals would be = 20/66 + 1/2 the frequency of Aa = 0.30 + 1/2 x 7/66 = 0.30 + 0.05 = 0.35 [where total number of individual = 66]

Hence, the frequency of the "a allele" would be = square under root of 0.35 = 0.59